OpenStudy (anonymous):

is 5/(2 times the squareroot (5x)) the same as square root of 5/(2 times the square root of x) ? i hope that made sense, Idk how to put the symbols

OpenStudy (anonymous):

\[\sqrt{5}/2\sqrt{x}..........and 5/2\sqrt{x}\]

OpenStudy (anonymous):

are you asking \[ {5 \over {2 \sqrt{5x}}} =? {5 \over {2\sqrt{x}}} \]

OpenStudy (anonymous):

or \[ {5 \over {2\sqrt{5x}}} =? \sqrt{5 \over {2 \sqrt{x}}} \]

OpenStudy (anonymous):

5/(2*sqrt(5x)) = sqrt(5)/(2sqrt(x)) Is true

OpenStudy (anonymous):

or \[ {5 \over {2\sqrt{5x}}} =? {\sqrt{5} \over {2 \sqrt{x}}} \]